The length of a rectangle is 2 inches more than its width.The area of the rectangle is 15 square inches. What are the length and width of the rectangle?
Answer by checkley77(12844)

L=W+2
AREA=LW
15=(W+2)W
15=W^2+2W
W^2+2W-15=0
(W+5)(W-3)=0
W-3=0
W=3 ANS.
L=3+2=5 ANS.
15=3*5
15=15